A TDMA system has a 2.5 MHz bandwidth with single channels of 25 kHz bandwidth each, with each frequency channel containing 4 TDMA conversation channels. If each user is assigned 1 transmit and 1 receive slot within each frequency channel, how many users can this system support?
Answer #1
Total Number of channel in system:- 2.5 * 106 / 25 * 103 = 100

Total Number of TDMA conversation channel= 4 * 100= 400

The Assign user gets 2 channels one for each direction of communication.

No of Users who can access this system:- 400/2 = 200

An suer

200